OpenStudy (anonymous):

find the common ratio for each of the following geometric sequence 128,96,72,54

OpenStudy (anonymous):

3/4

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle {\rm r}= a_{n+1} ~\div ~a_{n} }\) where, \(\color{#000000 }{ \displaystyle {\rm r} }\) - common ratio \(\color{#000000 }{ \displaystyle a_{n+1} }\) - (n+1)th term \(\color{#000000 }{ \displaystyle a_{n} }\) - (n)th term, (basically, some term before \(\color{black}{a_{n+1}}\))

OpenStudy (anonymous):

\[128(T_1),96(T_2),72(T_3),54(T_4),...\]\[r=\frac{ T_2 }{ T_1 }=\frac{ T_4 }{ T_3 }=...=\frac{ T_n }{ T_{n-1} }\]\[T_n=nth~term\]@banegirl@84

OpenStudy (anonymous):

@banegirl84